# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

### v3

@@ VI-B Open Source Dataset
The development of CP models relies on perception datasets gathered from various onboard sensors like radar, camera, and LiDAR. Utilizing these sensors for real-world perception tasks can be costly and time-consuming. A more economically viable alternative is to obtain cooperative data from simulation platforms featuring high-fidelity sensor modelling and perception capabilities. This subsection will outline several datasets relevant to CDA research, summarized in Table VIII.
@@
A direct approach to evaluating V2X CP involves collecting diverse real-world test scenarios. DAIR-V2X-C [273] is an autonomous driving dataset from real scenarios designed for research on V2I cooperation-related challenges. It comprises 71,254 LiDAR and 71,254 Camera frames captured at intersections equipped with infrastructure sensors. Expert annotators have meticulously labelled all frames with 3D annotations. While DAIR-V2X supports the perception task, it does not cover the essential motion prediction aspect.
V2X-Seq [274] offers a sequential V2I dataset, also collected from real-world scenarios, serving to develop and validate CP and motion forecasting algorithms. This dataset comprises two distinct sub-datasets: the sequential perception and trajectory forecasting datasets. The former includes over 15,000 frames from 95 scenarios, encompassing various data types such as infrastructure images, point clouds, vehicle-side images, point clouds, 3D detection/tracking annotations, and vector maps. The latter contains 210,000 scenarios, including 50,000 cooperative cases from 28 intersections, with a substantial data coverage of 672 hours.
For V2V scenarios, the V2V4Real dataset [275] was collected using two vehicles equipped with multi-modal sensors, driving through diverse scenarios in Columbus, Ohio, USA. In comparison to DAIR-V2X and V2X-Seq, V2V4Real covers a more extensive driving area and mileage, encompassing four road types: intersections, highway entrance ramps, highway straight roads, and city straight roads. This dataset comprises 40,000 RGB frames, 20,000 LiDAR frames, 240,000 annotated 3D bounding boxes for five object classes, and corresponding HDMaps for road topology prediction and semantic understanding of the BEV map. This substantial real-world dataset facilitates the training and testing of CP models for cooperative 3D object detection and tracking, as well as Sim2Real domain adaptation. It also offers comprehensive benchmarks involving eight state-of-the-art CP algorithms for these three tasks.
The V2V4Real dataset, with only two vehicles, lacks comprehensiveness for advancing research on cooperative applications. Addressing this limitation, LUCOOP [276] presents time-synchronized multi-modal data collected by three CAVs, each equipped with a broad sensor suite. This suite includes at least one LiDAR sensor, one GNSS antenna, and up to three IMUs. Additionally, Ultra-Wide-Band (UWB) sensors are mounted on each vehicle and placed statically along the trajectory to enable both V2V and V2X range measurements. The dataset encompasses sensor measurements in a multi-vehicle setup, spanning a trajectory of over 4 km and a time interval of more than 26 minutes. Overall, it includes more than 54,000 LiDAR frames, approximately 700,000 IMU measurements, over 2.5 hours of 10 Hz GNSS raw measurements, and 1 Hz data from a reference station [276]. The dataset also offers more than 6,000 total station measurements over a trajectory exceeding 1 km, as well as 1,874 V2V and 267 V2X UWB measurements [276]. Additionally, it provides 3D bounding box annotations for evaluating object detection methods and highly accurate ground truth poses for each vehicle throughout the measurement campaign [276].
The current practical utility of the LUCOOP dataset is limited by the absence of complex traffic scenarios. In contrast, the TUMTraf-V2X dataset [277] presents a comprehensive resource comprising 2,000 annotated point clouds and 5,000 annotated images collected from a variety of sensors. This dataset encompasses a total of 30,000 3D bounding boxes with associated track IDs, with a particular emphasis on challenging scenarios like near-misses, overtaking, U-turns, and traffic violations. Data from nine distinct sensors are included, capturing traffic scenes under diverse weather and lighting conditions. Additionally, the dataset includes rare occurrences, such as pedestrians crossing busy intersections against red lights. An open-source annotation tool and development kit are provided to facilitate the integration and evaluation of the dataset.
To address the absence of extensive datasets for real-world V2X communication, the study by [278] introduced a multi-modal dataset of ETSI ITS V2X messages gathered from public road traffic. This dataset includes over 230,000 ETSI ITS V2X messages from a diverse set of more than 1,800 vehicles and roadside units. The study provides details on the hardware setup and recording locations used to collect the V2AIX dataset, which encompasses urban and highway traffic sections in Germany. Furthermore, it provides insights into the current status of standardized V2X communication and offers a way to integrate these messages into the Robot Operating System (ROS) for automated driving applications.
@@
While realistic datasets allow CP models to generalize effectively to real-world driving situations, they are expensive, time-consuming, and limited in scenarios. A cost-effective alternative is to gather large-scale datasets in a high-fidelity simulator. The official version of the Cooperative Driving Dataset (CODD) [279], generated using the CARLA simulator, consists of 108 snippets. Each snippet contains 125 temporal frames of LiDAR point cloud data. Each frame in a snippet corresponds to a temporal data slice, including LiDAR data from all vehicles in the scenario, sensor absolute pose information, and ground-truth annotations for 3D bounding boxes of cars and pedestrians. CODD focuses on a single driving scenario. In contrast, the OPV2V dataset [280] incorporates camera images and LiDAR point clouds from connected vehicles across various scenes. Using OpenCDA and the CARLA simulator, OPV2V collects over 70 diverse scenes featuring varying numbers of connected vehicles. These scenes span eight towns within CARLA and a digital town, Culver City, Los Angeles. CODD and OPV2V provide ground truth for cooperative 3D object detection with V2V communication. However, neither dataset accounts for roadside infrastructure, limiting their ability to evaluate V2I perception.
The V2X-Sim dataset [281] serves as a synthetic V2X-assisted CP dataset in the field of autonomous driving. It is designed to stimulate research in multi-agent, multi-modality, and multi-task perception. The dataset was generated by SUMO to create numerically realistic traffic flows and CARLA to capture sensor streams from multiple vehicles at the same intersection. V2X-Sim offers well-synchronized recordings from roadside infrastructure and multiple vehicles at the intersection, along with comprehensive ground truth annotations to support various downstream tasks, including detection, tracking, and segmentation. However, it lacks noise simulation in its settings and only encompasses a single category of the road network.
@@
In contrast, the V2XSet dataset [198] explicitly accounts for real-world noise during V2X collaboration by leveraging CARLA and OpenCDA. It includes more than 11,447 frames divided into three subsets: training, validation, and testing, with ratios of 6,694, 1,920, and 2,833, respectively. The DOLPHINS dataset [282], generated using the CARLA traffic simulator, encompasses six distinct autonomous driving scenarios. It supports V2X communication and provides temporally synchronized sensor data from the ego vehicle, a cooperating vehicle, and a roadside unit. The dataset comprises 42,376 frames of sensor data and 292,549 3D bounding box annotations for cars and pedestrians.
@@
While the datasets above cover large-scale scenarios, they lack accident scenarios essential for comprehensive safety assessments of autonomous driving. Addressing this need, the DeepAccident dataset [283] was created through the CARLA simulator. The DeepAccident dataset comprises 57,000 frames and 285,000 samples, enabling the evaluation of motion and accident prediction capabilities. In each scenario, four vehicles and one infrastructure unit are established to capture sensor data, including multi-view RGB cameras and LiDAR. The dataset includes corresponding annotations for perception and motion prediction tasks. This dataset fills the gap by providing safety-critical scenarios absent in existing V2X datasets, offering diverse viewpoints for accident scenarios. Additionally, an end-to-end accident prediction task is introduced to forecast collision accidents, including their occurrence, timing, location, and the involved vehicles or pedestrians. The proposed V2XFormer serves as a benchmark for motion and accident prediction, as well as object detection, demonstrating superior performance in these predictive tasks compared to single-vehicle models.
VII Lessons Learned

### v4

@@ VI-B Open Source Dataset
The development of CP models relies on perception datasets gathered from various onboard sensors like radar, camera, and LiDAR. Utilizing these sensors for real-world perception tasks can be costly and time-consuming. A more economically viable alternative is to obtain cooperative data from simulation platforms featuring high-fidelity sensor modelling and perception capabilities. This subsection will outline several datasets relevant to CDA research, summarized in Table VII.
VI-B1 Real-world Dataset
@@
A direct approach to evaluating V2X CP involves collecting diverse real-world test scenarios. DAIR-V2X-C [292] is an autonomous driving dataset from real scenarios designed for research on V2I cooperation-related challenges. It comprises 38,845 LiDAR and 38,845 camera frames captured at intersections equipped with infrastructure sensors. Expert annotators have meticulously labelled all frames with 3D annotations. While DAIR-V2X supports the perception task, it does not cover the essential motion prediction aspect. V2X-Seq [293] offers a sequential V2I dataset, serving to develop and validate CP and motion forecasting algorithms. This dataset comprises two distinct sub-datasets: the sequential perception and trajectory forecasting datasets. The former includes over 15k frames from 95 scenarios, encompassing various data types such as infrastructure images, point clouds, vehicle-side images, point clouds, 3D detection/tracking annotations, and vector maps. The latter contains 80k infrastructure-view scenarios, 80k vehicle-view scenarios, and 50k cooperative-view scenarios captured from 28 intersections, with a substantial data coverage of 672 hours.
DAIR-V2X-C and V2X-Seq employ a single viewpoint utilizing a camera and Lidar pair to gather data from various intersections. Nevertheless, the intricacy of traffic scenarios often results in frequent occlusion of targets captured by the camera, thus significantly restricting the efficacy of single-viewpoint roadside sensors. In contrast, HoloVIC [294] was developed through the establishment of multiple holographic intersections equipped with diverse sensor configurations, including cameras, lidars, and fisheye cameras. It comprises over 100k synchronized frames and provides annotations for 3D bounding boxes, object IDs, and global trajectories. Furthermore, it delineates five distinct tasks to foster advancements in related research areas.
The V2V4Real dataset [295], collected from two CAVs with multi-modal sensors in Columbus, Ohio, USA, covers diverse driving scenarios such as intersections, highways, and city streets. It includes 40k RGB frames, 20k LiDAR frames, 240k annotated 3D bounding boxes, and HDMaps for road topology prediction, supporting cooperative 3D object detection and Sim2Real adaptation. However, its two-vehicle setup limits its comprehensiveness for broader cooperative applications. Addressing this, LUCOOP [296] provides data from three CAVs over a 4 km trajectory, including 54k LiDAR frames, 700k IMU measurements, 2.5 hours of GNSS data, UWB-based V2V and V2X measurements, and accurate ground truth poses. Although it offers a valuable resource for evaluating CP methods, it lacks complex traffic scenarios. In contrast, the TUMTraf-V2X [297] offers a more extensive collection, focusing on challenging traffic situations, varying weather conditions, and rare events, making it a more comprehensive tool for testing CP methods. It features 2k annotated point clouds, 5k images, and 30k 3D bounding boxes with track IDs.
The datasets mentioned above primarily support either V2I or V2V cooperation, but not both modes together. In contrast, the V2X-Real dataset [300] is the first real-world dataset designed specifically for V2X CP research. It includes 33K LiDAR frames, 171K camera images, and over 1.2 million annotated 3D bounding boxes across ten object categories, collected from two CAVs and two smart infrastructures equipped with multi-modal sensors. The dataset is organized into four sub-datasets based on the ego agent type and collaboration mode: Vehicle-Centric, Infrastructure-Centric, V2V, and I2I CP. V2X-Real is collected in congested urban environments with dense traffic and pedestrian flow, providing rich and challenging scenarios for V2X CP research.
All real-world datasets discussed so far have primarily focused on vehicle-centric cooperative tasks, such as V2V and V2I cooperation. However, roadside CP, particularly infrastructure-to-infrastructure (I2I) collaboration for comprehensive roadside coverage, remains unexplored. To address this, [298] introduced RCooper, the first real-world dataset for roadside CP, featuring over 50k images and 30k point clouds with 3D annotations across various weather, lighting, and traffic conditions. InScope [299] is another large-scale roadside CP dataset designed to tackle occlusion challenges in open traffic scenarios. It features multi-position LiDAR sensors deployed on infrastructure to broaden perception and complement vehicle-side systems, with 303 tracking trajectories and 187,787 3D bounding boxes captured over 20 days. InScope offers benchmarks for tasks like 3D object detection and multi-object tracking, providing comprehensive anti-occlusion algorithm evaluation. A new metric is also introduced to assess occlusion’s impact on detection performance.
VI-B2 Simulation Dataset
@@
While realistic datasets allow CP models to generalize effectively to real-world driving situations, they are expensive, time-consuming, and limited in scenarios. A cost-effective alternative is to gather large-scale datasets in a high-fidelity simulator. The Cooperative Driving Dataset (CODD) [301], generated using the CARLA simulator, comprises 108 snippets, each with 125 temporal frames of LiDAR point cloud data. Each frame contains LiDAR data from all vehicles in the scenario, sensor poses, and ground-truth annotations for 3D bounding boxes of cars and pedestrians. CODD focuses on a single driving scenario. In contrast, the OPV2V dataset [302] includes both camera images and LiDAR point clouds from connected vehicles across over 70 diverse scenes, spanning eight CARLA towns and Culver City, Los Angeles, using OpenCDA and CARLA. Both CODD and OPV2V provide ground truth for V2V cooperative 3D object detection, but neither includes roadside infrastructure, limiting their ability to evaluate V2I perception [308].
@@
The V2X-Sim dataset [303] is a synthetic V2X-assisted CP dataset for autonomous driving, aimed at advancing research in multi-agent, multi-modality, and multi-task perception. It was generated using SUMO for realistic traffic flows and CARLA for capturing sensor streams from multiple vehicles at intersections. V2X-Sim provides synchronized recordings from roadside infrastructure and vehicles, along with comprehensive ground truth annotations for tasks such as detection, tracking, and segmentation. However, it lacks noise simulation and only includes a single type of road network. In contrast, the V2XSet dataset [145] incorporates real-world noise during V2X collaboration using CARLA and OpenCDA, featuring over 11,447 frames divided into training, validation, and testing subsets. The DOLPHINS dataset [304], generated with the CARLA simulator, encompasses six autonomous driving scenarios with V2X communication and temporally synchronized sensor data from an ego vehicle, a cooperating vehicle, and a roadside unit. It contains 42,376 frames of sensor data and 292,549 3D bounding box annotations for cars and pedestrians.
@@
While the previously mentioned datasets cover large-scale scenarios, they lack accident scenarios crucial for a comprehensive safety assessment of autonomous driving. To address this gap, the DeepAccident dataset [305] was created using the CARLA simulator, featuring 57k frames and 285k annotated samples for evaluating motion and accident prediction capabilities. Data is captured by four vehicles and one infrastructure unit using multi-view RGB cameras and LiDAR, with corresponding annotations for perception and motion prediction tasks. This dataset provides safety-critical scenarios and introduces an end-to-end accident prediction task for forecasting collision details. V2XFormer, proposed as a benchmark, demonstrates superior performance in motion and accident prediction compared to single-vehicle models. However, DeepAccident lacks scenario diversity, realistic sensor models, environmental conditions, and vulnerable road users. The SCOPE dataset [306] addresses these limitations as the first synthetic multi-modal dataset for CP, incorporating realistic sensor models, physically accurate weather simulations, over 40 diverse scenarios with up to 24 collaborative agents, and passive traffic. It also provides benchmarks for 2D/3D object detection and semantic segmentation.
VII Performance Validation and Field Experiment
Evaluating CP systems requires a blend of simulations, offline testing, and real-world experiments. Simulations and offline validation allow for controlled performance assessments of V2X-based CP approaches, optimizing metrics to achieve optimal performance. In contrast, field experiments test CP adaptability in real-world conditions, revealing how systems handle diverse traffic and communication challenges. This section first compares the performance of various CP methods in offline testing and then explores real-world initiatives, including experimental deployments, pilot projects, and field tests, to assess CP readiness for practical deployment.
VII-A Performance Validation
In this subsection, performance validation is presented through a synthesis of experimental results from existing studies on CP methods for autonomous driving, using both real-world and simulated datasets, with 3D object detection used as an example downstream task. Average Precision (AP), defined as the area under the precision-recall curve, serves as the primary metric for 3D object detection, with APs at Intersection-over-Union (IoU) thresholds of 0.5 and 0.7 utilized to assess various models. Performance comparisons are presented across several datasets, including the real-world datasets DAIR-V2X [292] and V2V4Real [295], as well as simulation-based datasets OPV2V [302], V2X-Sim [303], and V2XSet [145].
VII-A1 Detection Performance Comparison
To ensure fairness and generalizability, all reported results reflect the best outcomes from experimental evaluations of existing methods, with only those methods assessed on at least two datasets included. Table VIII displays the AP values at IoU thresholds of 0.5 and 0.7 across five datasets under default settings, highlighting the highest AP values and indicating the source for each result. The observed performance differences across methods and datasets indicate that certain algorithms are more effective under specific environmental or data conditions.
Table VIII results indicate that DI-V2X achieves the highest AP@0.5 on DAIR-V2X at 0.7882, while DiscoNet leads on V2V4Real with an AP@0.5 of 0.7360. For AP@0.7 on V2V4Real, both FPV-RCNN and MACP achieve top performance at 0.4790, though MACP scores slightly lower than FPV-RCNN with an AP@0.5 of 0.6760. Among simulation-based datasets, MKD-Cooper performs exceptionally on V2X-Sim, with an AP@0.5 of 0.9300 and AP@0.7 of 0.8520, and achieves a high AP@0.7 of 0.9240 on OPV2V, positioning it as a leading method in simulated environments. V2X-PC demonstrates robust performance across both real and simulated datasets, achieving an AP@0.5 of 0.7689 and AP@0.7 of 0.6939 on DAIR-V2X and even higher scores on V2XSet with an AP@0.5 of 0.9283 and AP@0.7 of 0.8955. These findings suggest that V2X-PC excels across diverse datasets, indicating strong adaptability to various conditions.
TABLE VIII: 3D Object Detection Performance Comparison On DAIR-V2X, V2V4Real, OPV2V, V2XSet, and V2X-Sim datasets, with AP@0.5 and AP@0.7.
Method Year Real-world Datasets Simulation-based Datasets
DAIR-V2X V2V4Real OPV2V V2XSet V2X-Sim
AP@0.5/0.70.50.70.5/0.70.5 / 0.7 Source AP@0.5/0.70.50.70.5/0.70.5 / 0.7 Source AP@0.5/0.70.50.70.5/0.70.5 / 0.7 Source AP@0.5/0.70.50.70.5/0.70.5 / 0.7 Source AP@0.5/0.70.50.70.5/0.70.5 / 0.7 Source
F-Cooper [203] 2019 $0.7370/0.5600$ [193] $0.6930/0.4320$ [193] $0.9070/0.8100$ [263] $0.8400/0.6800$ [145] $0.7150/0.5470$ [218]
V2VNet [140] 2020 $0.7097/0.4763$ [229] $0.6470/0.3360$ [216] $0.9350/0.7400$ [190] $0.8710/0.6460$ [191] $0.8850/0.6960$ [218]
When2com [116] 2020 $0.5188/0.3705$ [129] - - $0.7785/0.6240$ [142] $0.7016/0.5372$ [196] $0.5206/0.4421$ [132]
Who2com [115] 2020 - - - - $0.5536/0.2339$ [132] - - $0.4977/0.4230$ [132]
V2VNetrobust [189] 2021 $0.6610/0.4860$ [193] $0.5500/0.3090$ [193] $0.9420/0.8540$ [193] - - $0.8400/0.7540$ [190]
DiscoNet [144] 2021 $0.7370/0.5840$ [193] 0.7360$/0.4660$ [193] ${0.9160}/0.7910$ [190] $0.9078/0.8381$ [136] $0.8180/0.5370$ [218]
OPV2V [302] 2022 $0.7330/0.5530$ [190] $0.6450/0.3430$ [216] $0.9430/0.8270$ [190] $0.9188/0.8475$ [136] $0.8870/0.7340$ [218]
Where2comm [139] 2022 $0.7520/0.5880$ [193] $0.7040/0.4690$ [193] $0.9440/0.8550$ [193] $0.9130/0.8530$ [232] $0.6205/0.5459$ [132]
V2X-Vit [145] 2022 $0.7398/0.6150$ [136] ${0.6800}/0.3910$ [193] $0.9460/0.8560$ [190] $0.9100/0.8030$ [191] $0.9010/0.7950$ [218]
CoBEVT [214] 2022 $0.6390/0.5167$ [136] $0.6650/0.3600$ [216] $0.9140/0.8620$ [262] $0.9033/0.8269$ [136] - -
FPV-RCNN [131] 2022 $0.6550/0.5050$ [190] 0.7010/0.7010/0.7010 /0.4790 [193] $0.8580/0.8400$ [190] $0.8650/0.5630$ [191] $0.8700/0.8380$ [190]
UMC [132] 2023 - - - - $0.6190/0.2450$ [132] - - $0.6780/0.6001$ [132]
What2comm [150] 2023 $0.6081/0.4463$ [150] - - $0.8683/{0.7361}$ [150] $0.8459/0.6686$ [150] - -
SCOPE [196] 2023 $0.6518/0.4989$ [196] - - $0.8971/0.8062$ [196] $0.8752/0.7505$ [196] - -
BM2CP [129] 2023 $0.6403/0.4899$ [129] - - $0.8372/0.6317$ [129] - - - -
CoAlign [190] 2023 $0.7460/0.6040$ [190] $0.7090/0.4170$ [193] 0.9660$/0.9120$ [190] $0.9190/0.8050$ [191] $0.8580/0.7650$ [190]
V2VFormer [216] 2024 - - 0.6540///0.3610 [216] $0.9170/0.8760$ [216] - - 0.6150///0.5500 [216]
CoBEVGlue [193] 2024 0.7400///0.5820 [193] 0.7020///0.4310 [193] 0.9580///0.9090 [193] - - - -
IFTR [134] 2024 0.2058///0.0665 [134] - - 0.8556///0.6604 [134] 0.7173///0.4967 [134] - -
V2X-M2C [232] 2024 - - - - 0.8870///0.8280 [232] 0.9220///0.8600 [232] - -
CPSC [142] 2024 0.6542///0.4784 [142] - - 0.8911///0.7797 [142] - - - -
V2X-INCOP [236] 2024 0.6548///0.5526 [236] - - 0.9256///0.8318 [236] - - - -
RoCo[191] 2024 0.7630///0.6200 [191] - - - - 0.9190///0.8050 [191] - -
MACP [228] 2024 - - 0.6760///0.4790 [228] 0.9370/0.9030 [228] - - - -
DI-V2X [229] 2024 0.7882///0.6616 [229] - - - - 0.9270///0.8270 [229] - -
MKD-Cooper [218] 2024 - - - - 0.9650/0.9650/0.9650 /0.9240 [218] - - 0.9300///0.8520 [218]
V2X-PC [136] 2024 0.7689///0.6939 [136] - - - - 0.9283///0.8955 [136] - -
VII-A2 Robustness to Time Delay
Time delay introduced by V2X communication presents a significant challenge in CP, leading to data asynchronization between ego and other agents during the data fusion process. To illustrate model robustness to time latency, we present performance across delays ranging from 0 to 0.4 seconds, as shown in Fig. 6, with results referenced from [150] and [217].
As shown in Fig. 6, detection performance declines for all methods as time delay increases. V2VFormer++ maintains relatively high AP scores, demonstrating the greatest resilience to latency. CoBEVT and What2comm also show robust performance, though to a lesser extent. Where2com and V2X-ViT exhibit moderate robustness, with gradual AP declines indicating some sensitivity to delay. In contrast, DiscoNet and V2VNet experience sharper AP declines, reflecting higher latency sensitivity. When2com begins with lower AP scores and degrades significantly as delay increases, highlighting its vulnerability to latency. In summary, V2VFormer++ proves the most delay-tolerant, while When2com is most affected by latency in the OPV2V setting, limiting its effectiveness under substantial transmission delays.
Figure 6: The impact of time delay on the AP@0.7 performance metric for the OPV2V dataset. The results are referenced from [150] and [217].
VII-A3 Robustness to Pose Errors
Collaborative agents rely on precise pose data from others to accurately transform the coordinates of the data in the received messages. However, even with advanced localization technologies such as GPS, pose errors remain unavoidable, requiring collaborative approaches to be resilient to these inaccuracies. Following [190], Gaussian noise $\mathcal{N}(0,\sigma_{t})$ is added to $x,y$, and $\mathcal{N}(0,\sigma_{r})$ to $\theta$, to simulate pose errors, where $x,y,\theta$ represent the 2D center and yaw angle of accurate global poses. Fig. 7 shows the detection performance in terms of AP@0.7 for ten different methods on the DAIR-V2X dataset, evaluated at four levels of pose error $(0.0m/0.0^{\circ},0.2m/0.2^{\circ},0.4m/0.4^{\circ},0.6m/0.6^{\circ})$, with results sourced from [190] and [191].
Among these methods, RoCo maintains the highest AP scores across noise levels, demonstrating strong resilience to pose errors. In contrast, V2VNet shows the lowest AP with substantial declines at higher noise levels, while its robust variant, V2VNetrobust, outperforms the standard version, indicating improved robustness to pose noise. CoAlign and CoBEVFlow perform well but yield slightly lower AP than RoCo. F-Cooper, DiscoNet, OPV2V, and V2X-ViT display stable performance across noise levels, suggesting greater resilience to pose errors. FPV-RCNN, despite moderate detection accuracy, shows notable AP declines as noise increases, indicating higher sensitivity to positional errors. Overall, increasing pose errors reduce AP@0.7 across all methods, underscoring their sensitivity to positional inaccuracies.
Figure 7: Detection performance on DAIR-V2X dataset with pose errors at four different noise levels. The pose noises follow Gaussian distribution $\mathcal{N}(0,\sigma_{t})$ on $x,y$ and $\mathcal{N}(0,\sigma_{r})$ on $\theta$, where $x,y,\theta$ are 2D centers and yaw angle of accurate global poses. We present AP@0.7 at noise level $\sigma_{t}/\sigma_{r}=0.0m/0.0^{\circ},0.2m/0.2^{\circ},0.4m/0.4^{\circ},0.6m/0.6^{\circ}$. The results are referenced from [190] and [191].
VII-B Field Experiments and Pilot Projects
Field experiments of CP systems have progressed through a range of global pilot projects, demonstrating CP’s adaptability and effectiveness under specific conditions. These initiatives showcase CP’s potential to improve traffic efficiency, safety, and the functionality of CAVs within controlled environments, advancing research and development in intelligent transportation systems.
The Cooperative Adaptive Cruise Control (CACC) project carried out in the U.S. leverages DSRC-based communication to enable real-time sharing of speed and positioning data between vehicles, allowing for autonomous maintenance of safe inter-vehicle distances [309]. The project aims to enhance Adaptive Cruise Control (ACC) by dynamically coordinating vehicle spacing within a string, ultimately improving traffic flow. Test cases for this project include simultaneous passing of vehicles on adjacent lanes, driving in a string with induced oscillations to assess string stability, handling communication interruptions in CACC mode, and evaluating the effects of positional inaccuracies in preceding or adjacent vehicles.
The Autonomous Intersection Management (AIM) project carried out in the U.S. explores autonomous traffic management at intersections without traditional traffic signals, using a scalable multi-agent framework suitable for urban areas [310]. Recognizing the coexistence of autonomous and human-driven vehicles, AIM introduced the “Hybrid AIM” protocol to facilitate safer, more efficient integration at intersections. The project employs augmented reality to create a mixed-reality setup, simulating virtual traffic on real roads, allowing for comprehensive testing of CP-enabled systems in dynamic and complex scenarios. This project offers a valuable template for urban areas aiming to reduce congestion through CP, highlighting how CP can enhance intersection management by enabling seamless coordination among diverse vehicle types.
The Ko-PER Project, carried out in Germany, focuses on road safety, using distributed sensor networks to build a comprehensive view of local traffic [311]. By integrating data from vehicle-mounted and infrastructure sensors, Ko-PER enables the detection of occluded traffic participants and the evaluation of collision risk [312]. A dataset has been produced from this project, containing laser scanner and video data from public intersections, and supports ongoing research on intersection perception and cooperative warning systems [313].
The Virginia Connected Corridors (VCC) project in the U.S. is led by the Virginia Department of Transportation and the Virginia Tech Transportation Institute [314]. It serves as a testbed for Connected and Autonomous Vehicle (CAV) applications. Since its launch in 2016, a fleet of 50 highly instrumented light vehicles and 60 Roadside Units (RSUs) have been deployed within the Northern Virginia Test Bed for application testing. Each vehicle is equipped to communicate via DSRC with the surrounding connected infrastructure. When the vehicles are outside the range of a DSRC-equipped RSU, they can use cellular technology for communication. A centralized cloud system is employed to manage the message traffic between connected vehicles. The vehicles can connect to the cloud through either DSRC or cellular networks. It has been noted that cellular communications have a longer latency, ranging from 1.5 to 3.5 seconds, compared to DSRC and are therefore used for applications where timely message receipt is not critical. The aim of this project is to provide an open development platform, allowing third-party developers to test and deploy applications that enhance roadway safety and mobility.
The iMOVE Cooperative Perception Project, led by the Australian Centre for Field Robotics at the University of Sydney in collaboration with Cohda Wireless, has demonstrated advancements in pedestrian detection around corners through the use of CPMs and Dedicated Short-Range Communication (DSRC) technology [315]. Notably, this project illustrates how CAVs can autonomously and safely interact with pedestrians, both walking and running, relying exclusively on CP information provided by an intelligent roadside unit (RSU) [315]. This project represents one of the pioneering demonstrations of urban vehicle automation that relies solely on CP data for safe and efficient operation in complex environments.
These projects demonstrate the versatility of CP across various traffic scenarios, including highway platooning, intersection management, and pedestrian detection in urban areas. They underscore CP’s important role in addressing perception gaps caused by occlusions or limited sensor range, ultimately leading to more reliable and robust autonomous driving systems. Additionally, the challenges faced during these projects will inform the development of next-generation CP algorithms and infrastructure to create safer and more efficient autonomous transportation ecosystems.
VIII Lessons Learned